# GeneMAN: Generalizable Single-Image 3D Human Reconstruction from Multi-Source Human Data

### v1

@@ Abstract
Figure 1: GeneMAN is a generalizable framework for single-view-to-3D human reconstruction, built on a collection of multi-source human data. Given a single in-the-wild image of a person, GeneMAN could reconstruct a high-quality 3D human model, regardless of its clothing, pose, or body proportions (e.g., a full-body, a half-body, or a close-up shot) in the given image. The project page of GeneMAN is: https://roooooz.github.io/GeneMAN/.
@@ 1 Introduction
In this paper, we propose GeneMAN, a Generalizable image-to-3D huMAN reconstruction framework for high-fidelity 3D human reconstruction from a single in-the-wild image. To enhance generalizability, we first collect a comprehensive, multi-source training dataset of high-quality, multi-modal human data, including 3D scans, multi-view videos, single images, and our augmented human data. Based on the multi-source human data collection, human-specific prior models, including a text-to-image diffusion model as the 2D prior and a view-conditioned diffusion model as the 3D prior, have been trained, which could provide more generalizable human priors compared to traditional human parametric models. Leveraging the pretrained human prior models, GeneMAN generates 3D human models through two primary stages. 1) 3D Geometry Reconstruction. To create precise and intricate 3D human geometry, GeneMAN employs an Initialization-&\&&-Sculpting strategy, which initially predicts a coarse human geometry using NeRF [36], followed by a sculpting process for adding geometry details. 2) 3D Texture Generation. Utilizing the refined human geometry, GeneMAN implements a multi-space texture refinement pipeline to produce high-fidelity, consistent 3D textures. Initially, coarse textures are created through multi-view texturing and iteratively refined in the latent space. Subsequently, detailed 3D textures are achieved via pixel space texture refinement, optimizing the UV maps with a 2D prior-based ControlNet [67]. Extensive experimental results demonstrate that GeneMAN surpasses existing state-of-the-art (SoTA) methods, showcasing strong generalization capability and high generation quality. We would like to highlight that GeneMAN could faithfully reconstruct 3D human models with diverse clothing, poses, and different personal belongings, given a single in-the-wild image with varying body proportions.
Our contributions are summarized as follows:
  • We propose a generalizable 3D human reconstruction framework called GeneMAN, building upon human-specific prior models trained on a multi-source human data collection.
@@ 2 Related Work
Template-based 3D Human Reconstruction. Template-based approaches rely on human parametric models [30, 39, 64] to guide 3D reconstruction [72, 62, 63, 19, 11, 70, 71, 13]. For instance, PaMIR [72] and ICON [62] leverage SMPL-derived features to facilitate implicit surface regression. By contrast, ECON [63] employs explicit representations, predicting normals and depths to generate 2.5D surfaces. GTA [70] employs transformers with learnable embeddings to translate image features into 3D triplane features. Recent methods [18, 13, 71] exploit the generative capabilities of pretrained diffusion models [43] for 3D reconstruction. SiTH [13] finetunes an image-conditioned diffusion model to hallucinate back views for 3D mesh reconstruction. SIFU [71] and TeCH [18] utilize diffusion-based 3D geometry and texture refinement to enhance the reconstruction quality. However, template-based methods are highly reliant on precise human parametric models, which are frequently unattainable for in-the-wild human images.
Template-Free 3D Human Reconstruction. Without using parameterized human models, PIFu [46] introduces a novel approach by extracting pixel-aligned features to build neural fields. Likewise, PIFuHD [47] improves this with high-resolution normal guidance, and PHOHRUM [4] enhances the pipeline by jointly predicting 3D geometry, surface albedo, and shading. Recently, HumanSGD [3] leverages pretrained 2D diffusion models [43] as a human appearance prior. LRM-Human [60] utilizes a considerable amount of human data to facilitate template-free human reconstruction. Nonetheless, they still encounter challenges, such as unrealistic and inconsistent textures, and inferior geometric details, largely due to insufficient priors and overly simplistic model designs.
@@ 3 Preliminaries
Diffusion Model. Diffusion models [14, 54, 55] learn to gradually transform samples from a tractable noise distribution towards a data distribution. These models comprise a forward process, $\{q_{t}\}_{t\in[0,1]}$, which incrementally adds random noise to a data point $\bm{x}_{0}\sim q_{0}(\bm{x}_{0})$, and a reverse process, $\{p_{t}\}_{t\in[0,1]}$, which gradually recovers clean data from noise. The forward process is defined by a conditional Gaussian distribution: $q_{t}(\bm{x}_{t}|\bm{x}_{0}):=\mathcal{N}(\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I})$, where $\alpha_{t},\sigma_{t}>0$ are time-dependent coefficients. The reverse process is defined by denoising from $p_{1}(\bm{x}_{1})\coloneqq\mathcal{N}(\bm{0},\bm{I})$ with a parameterized noise prediction network $\epsilon_{\phi}(\bm{x}_{t},t)$ to predict the noise added to a clean data $\bm{x}_{0}$, which is trained by minimizing
@@
Recently, Zero-1-to-3 [28] finetunes a variant of Stable Diffusion [43] model on the Objaverse [8] dataset, resulting in a view-conditioned diffusion model $\phi$, which could synthesize novel view images at any arbitrary viewpoint $\bm{c}$ given a reference image $\bm{\hat{I}}$. Hence, $\phi$ offers a strong 3D prior, based on which could derive a 3D-aware SDS loss $\mathcal{L}_{\textnormal{SDS}}^{3D}$. Its gradient $\nabla_{\theta}\mathcal{L}_{\textnormal{SDS}}^{3D}$ is formulated as:
@@
4 Method
In this section, we propose GeneMAN, a generalizable image-to-3D human pipeline aimed at reconstructing high-quality 3D humans from in-the-wild images. By leveraging GeneMAN prior models, which provide human-specific geometric and texture priors, we achieve high-fidelity 3D human reconstruction with varying body proportions, diverse poses, clothing, and personal belongings. The success of GeneMAN prior models is closely tied to the multi-source human dataset we constructed. We first introduce the construction of a multi-source human dataset and GeneMAN prior models in Sec. 4.1 and Sec. 4.2, respectively. Then, we illustrate the details of our GeneMAN framework, including the Geometry Initialization &\&& Sculpting (Sec. 4.3), as well as the Multi-Space Texture Refinement (Sec. 4.4). The overview of GeneMAN framework is shown in Fig. 2.
@@ 4.1 Multi-Source Human Dataset
Our 3D scanned human data is aggregated from the commercial dataset RenderPeople [1], alongside several open-source datasets: CustomHumans [12], HuMMan [6], THuman2.0 [65], THuman3.0 [56] and X-Humans [50]. Additionally, we enrich the dataset by integrating human-specific data filtered from Objaverse [8]. For multi-view human videos, we leverage datasets such as DNA-Rendering [7], ZJU-Mocap [40], AIST++ [22], Neural Actor [27] and Actors-HQ [20]. In terms of 2D human imagery, we select data from DeepFashion [29] and LAION-5B [48] to ensure comprehensive coverage of diverse human appearances. Furthermore, we employ data augmentation strategies to synthesize additional data. Specifically, we utilize ControlNet-based [67] image synthesis to generate multi-view human data with diverse clothing options through various prompts, enabling the creation of human images featuring different garments. To account for varying body proportions, we apply image cropping to preprocessed multi-view human renderings, expanding our dataset to include instances with diverse body proportions. In total, our dataset comprises over $50K$ multi-view instances. Further details of the multi-source human dataset are provided in the supplementary materials.
@@ 4.3 Geometry Initialization &\&& Sculpting
where $g(\theta;\hat{c})$ represents the rendered image at the reference viewpoint $\hat{c}$, $\theta$ denotes the NeRF parameters to be optimized and $\odot$ indicates the Hadamard product. The foreground region mask is denoted as $\bm{\hat{m}}$, and $g_{m}(\theta;\hat{c})$ renders the silhouette. In addition to supervising the RGB image and mask, we also exploit the geometry prior inferred from the reference image, specifically depth and normal information. The reference depth $\bm{\hat{I}_{d}}$ and normal $\bm{\hat{I}_{n}}$ are derived using the off-the-shelf human foundation model Sapiens [21]. To enforce consistency between the rendered and reference values, we impose depth and normal losses, $\mathcal{L}_{\textnormal{depth}}$ and $\mathcal{L}_{\textnormal{normal}}$. The normal loss is calculated as the mean squared error (MSE) between the reference normal $\bm{\hat{I}_{n}}$ and $\bm{I_{n}}$, i.e., $\mathcal{L}_{\textnormal{normal}}=||\bm{\hat{I}_{n}}-\bm{I_{n}}||_{2}$. For the depth loss, we employ the normalized negative Person correlation as follows:
@@
To further refine the geometric details, we adopt DMTet [51], a hybrid SDF-Mesh representation that enables memory-efficient, high-resolution 3D shape generation. The trained NeRF [36] is converted into a mesh, which serves as the geometric initialization for DMTet optimization. During optimization, we compute the reconstruction loss by applying both the MSE loss and the perceptual loss [53] between the rendered normal from the reference view and the reference normal $\bm{\hat{I}_{n}}$. Inspired by HumanNorm [17], we leverage its pretrained normal-adapted and depth-adapted diffusion models for novel view guidance to craft high-quality geometry. Furthermore, the SDF loss $\mathcal{L}_{\textnormal{sdf}}$ is incorporated to regularize the geometry, preventing significant deviation from the initial mesh and avoiding improper shapes. Through geometry initialization and sculpting, we achieve high-fidelity 3D human geometry with plausible poses and intricate geometric details.
@@ 4.4 Multi-Space Texture Refinement
where $\lambda^{c}_{ref}$ denotes the weighting parameter of the reference loss, and $\mathcal{L}^{c}_{\textnormal{guid}}$ represents the guidance loss in coarse texture stage, as defined in Eq. 6.
Multi-View Training. Multi-view training strategies have proven effective in enhancing multi-view consistency in several methods [28, 52]. However, these approaches often require the retraining of new diffusion models. In contrast, we employ a training-free approach by adding identical Gaussian noise to a batch of images rendered from different views. These images are then concatenated into a single image for inference, ensuring that the resulting multi-view images exhibit consistent textures and colors.
The coarse texture stage prioritizes learning consistent and plausible texture but still lacks realism. To further enhance texture realism, we propose a pixel space optimization by minimizing the distance between the rendered multi-view images and their refined counterparts, thereby improving the UV texture map. Akin to DreamGaussian [58], we adopt the image-to-image synthesis framework of SDEdit [35] to generate the refined images. Specifically, we render the coarse image $\bm{I_{\textnormal{corase}}}$ from an arbitrary camera view $c$, perturb it with random Gaussian noise, and apply a multi-step denoising process using the GeneMAN 2D prior based ControlNet [67] $\phi_{2d}(\cdot)$ to obtain the refined image $\bm{I_{\textnormal{fine}}}$:
@@
where $\epsilon(t_{\textnormal{start}})$ represents random noise at timestep $t_{\textnormal{start}}$, and $e$ denotes the conditioned text embeddings. We employ MSE loss for pixel-wise reconstruction, supplemented by LPIPS [68] loss to enhance texture details and style similarity. The total loss for pixel space optimization $\mathcal{L}_{\textnormal{fine}}$ is:
@@ 5.1 Implementation Details
Training Details. Our GeneMAN framework is built upon the open-source project ThreeStudio [10]. During the geometry stage, we progressively increase the resolution of NeRF [36] from $256$ to $384$ over $5,000$ steps. We then convert it to an explicit mesh, which serves as the geometry initialization for DMTet [51] at a high resolution of $512$. We subsequently optimize DMTet for $10,000$ steps to sculpt fine-grained geometric details. In the texture stage, we perform an initial coarse texture optimization over $10,000$ steps, followed by a refinement of the UV map texture for additional $1,000$ steps. The total optimization process takes approximately two hours. Additional training details are provided in the supplementary materials.
Testing Details. For both qualitative and quantitative evaluation, we randomly select $50$ samples from CAPE [33] and the Internet. The samples sourced from the internet include challenging scenarios, such as humans with varying body proportions (head shots, half-body shots, full-body shots), as well as humans in diverse poses, clothing, and with personal belongings. We compare our model with state-of-the-art image-to-3D human reconstruction methods, including PIFu [46], GTA [69], TeCH [18] and SiTH [13]. For each method, we render 120 viewpoints across 360 degrees of the reconstructed results for evaluation. We compute PSNR and LPIPS [68] between the reference view and the rendered front view to evaluate reconstruction quality. Additionally, CLIP similarity [42] between the reference view and $119$ novel views is measured to assess multi-view consistency.
Table 1: Quantitative Comparison with State-of-the-art Methods on CAPE [33]. The best results are highlighted in bold. The second-place results are underlined.
[…]
PIFu 26.912 0.028 0.764
GTA 30.376 0.019 0.785
TeCH 27.601 0.025 0.826
SiTH 21.992 0.048 0.815
GeneMAN (Ours) 28.490 0.015 0.838
Table 2: Quantitative Comparison with State-of-the-art Methods on in-the-wild images. The best results are highlighted in bold. The second-place results are underlined.
[…]
PIFu 26.968 0.035 0.594
GTA 25.060 0.064 0.568
TeCH 25.740 0.053 0.713
SiTH 20.412 0.129 0.608
GeneMAN (Ours) 32.238 0.013 0.730

### v2

@@ Abstract
Figure 1: GeneMAN is a generalizable framework for single-view-to-3D human reconstruction, built on a collection of multi-source human data. Given a single in-the-wild image of a person, GeneMAN could reconstruct a high-quality 3D human model, regardless of its clothing, pose, or body proportions (e.g., a full-body, a half-body, or a close-up shot) in the given image. The anonymous project page of GeneMAN is: https://roooooz.github.io/GeneMAN/.
@@ 1 Introduction
In this paper, we propose GeneMAN, a Generalizable image-to-3D huMAN reconstruction framework for high-fidelity 3D human reconstruction from a single in-the-wild image. To enhance generalizability, we first collect a comprehensive, multi-source training dataset of high-quality, multi-modal human data, including 3D scans, multi-view videos, single images, and our augmented human data. Based on the multi-source human data collection, human-specific prior models, including a text-to-image diffusion model as the 2D prior and a view-conditioned diffusion model as the 3D prior, have been trained, which could provide more generalizable human priors compared to traditional human parametric models. Leveraging the pretrained human prior models, GeneMAN generates 3D human models through two primary stages. 1) 3D Geometry Reconstruction. To create precise and intricate 3D human geometry, GeneMAN employs an Initialization-&\&&-Sculpting strategy, which initially predicts a coarse human geometry using NeRF [36], followed by a sculpting process for adding geometry details. 2) 3D Texture Generation. Utilizing the refined human geometry, GeneMAN implements a multi-space texture refinement pipeline to produce high-fidelity, consistent 3D textures. Initially, coarse textures are created through multi-view texturing and iteratively refined in the latent space. Subsequently, detailed 3D textures are achieved via pixel space texture refinement, optimizing the UV maps with a 2D prior-based ControlNet [68]. Extensive experiments demonstrate that GeneMAN surpasses existing state-of-the-art (SoTA) methods, showcasing strong generalization capability and high generation quality. We would like to highlight that GeneMAN could faithfully reconstruct 3D human models with diverse clothing, complex poses, and different personal belongings, given a single in-the-wild image with varying body proportions. Our contributions are summarized as follows:
  • We introduce GeneMAN, a generalizable 3D human reconstruction framework built on human-specific prior models trained on collected multi-source data. Our framework enables high-quality 3D human reconstruction regardless of its clothing, pose, or body proportions.
@@ 2 Related Work
Template-based 3D Human Reconstruction. Template-based approaches rely on human parametric models [30, 39, 65] to guide 3D reconstruction [73, 63, 64, 19, 11, 71, 72, 13]. PaMIR [73] and ICON [63] leverage SMPL-derived features to facilitate implicit surface regression. By contrast, ECON [64] employs explicit representations, predicting normals and depths to generate 2.5D surfaces. GTA [71] utilizes transformers with learnable embeddings to translate image features into 3D triplane features. Several methods [18, 13, 72] exploit the generative capabilities of pretrained diffusion models [44] for 3D human reconstruction. SiTH [13] finetunes an image-conditioned diffusion model to hallucinate back views, while SIFU [72] and TeCH [18] utilize diffusion-based geometry and texture refinement to improve reconstruction quality. Recently, IDOL [74] predicts Gaussian-based representation of 3D human to enable fast reconstruction. However, template-based methods are highly reliant on precise human parametric models, which are frequently unattainable for in-the-wild human images.
Template-Free 3D Human Reconstruction. Without parameterized human models, PIFu [47] introduces a novel approach by extracting pixel-aligned features to build neural fields. PIFuHD [48] improves this with high-resolution normal guidance, while PHOHRUM [4] jointly predicts 3D geometry, surface albedo, and shading. Recently, HumanSGD [3] leverages pretrained 2D diffusion models [44] as a human appearance prior. HumanLRM [61] trains a diffusion-guided feed-forward model to predict the implicit field of a human without template priors. Nonetheless, they still encounter challenges, such as unrealistic and inconsistent textures, and inferior geometric details, largely due to insufficient priors and overly simplistic model designs.
@@ 3 Preliminaries
Diffusion Model. Diffusion models [14, 55, 56] learn to transform samples from a tractable noise distribution towards a data distribution. They comprise a forward process $\{q_{t}\}_{t\in[0,1]}$, which progressively adds random noise to a data point $\bm{x}_{0}\sim q_{0}(\bm{x}_{0})$, and a reverse process $\{p_{t}\}_{t\in[0,1]}$, which gradually recovers clean data from noise. The forward process is defined by a conditional Gaussian distribution: $q_{t}(\bm{x}_{t}|\bm{x}_{0}):=\mathcal{N}(\alpha_{t}\bm{x}_{0},\sigma_{t}^{2}\bm{I})$, where $\alpha_{t},\sigma_{t}>0$ are time-dependent coefficients. The reverse process is defined by denoising from $p_{1}(\bm{x}_{1})\coloneqq\mathcal{N}(\bm{0},\bm{I})$ with a parameterized noise prediction network $\epsilon_{\phi}(\bm{x}_{t},t)$ to predict the noise added to a clean data $\bm{x}_{0}$, which is trained by minimizing
@@
A pivotal work, Zero-1-to-3 [28] finetunes a variant of Stable Diffusion [44] model on the Objaverse [8] dataset, resulting in a view-conditioned diffusion model $\phi$, which could synthesize novel view images at any arbitrary viewpoint $\bm{c}$ given a reference image $\bm{\hat{I}}$. Hence, $\phi$ offers a strong 3D prior, based on which could derive a 3D-aware SDS loss $\mathcal{L}_{\textnormal{SDS}}^{3D}$. Its gradient $\nabla_{\theta}\mathcal{L}_{\textnormal{SDS}}^{3D}$ is formulated as:
@@
4 Method
In this section, we propose GeneMAN, a generalizable image-to-3D human pipeline aimed for reconstructing high-quality 3D humans from in-the-wild images. By leveraging GeneMAN prior models, which provide human-specific geometric and texture priors, we achieve high-fidelity 3D human reconstruction with varying body proportions, diverse poses, clothing, and personal belongings. The success of GeneMAN prior models is closely tied to the multi-source human dataset we constructed. We first introduce the construction of a multi-source human dataset and GeneMAN prior models in Sec. 4.1 and Sec. 4.2, respectively. Then, we illustrate the details of our GeneMAN framework, including the Geometry Initialization &\&& Sculpting (Sec. 4.3), as well as the Multi-Space Texture Refinement (Sec. 4.4). The overview of GeneMAN framework is shown in Fig. 2.
@@ 4.1 Multi-Source Human Dataset
Our 3D scanned human data is aggregated from the commercial dataset RenderPeople [1], alongside several open-source datasets: CustomHumans [12], HuMMan [6], THuman2.0 [66], THuman3.0 [57] and X-Humans [51]. Additionally, we enrich the dataset by integrating human-specific data filtered from Objaverse [8]. For multi-view human videos, we leverage datasets such as DNA-Rendering [7], ZJU-Mocap [40], AIST++ [22], Neural Actor [27] and Actors-HQ [20]. In terms of 2D human imagery, we select data from DeepFashion [29] and LAION-5B [49] to ensure comprehensive coverage of diverse human appearances. Furthermore, we employ data augmentation strategies to synthesize additional data. Specifically, we utilize ControlNet-based [68] image synthesis to generate multi-view human data with diverse clothing options through various prompts, enabling the creation of human images featuring different garments. To account for varying body proportions, we apply image cropping to preprocessed multi-view human renderings, expanding our dataset to include instances with diverse body proportions. In total, our dataset comprises over $50K$ multi-view instances. Further details of the dataset are provided in the supplementary materials.
@@ 4.3 Geometry Initialization &\&& Sculpting
where $g(\theta;\hat{c})$ represents the rendered image at the reference viewpoint $\hat{c}$, $\theta$ denotes the NeRF parameters and $\odot$ indicates the Hadamard product. The foreground mask is denoted as $\bm{\hat{m}}$, and $g_{m}(\theta;\hat{c})$ renders the silhouette. In addition to supervising the RGB image and mask, we utilize the geometry prior inferred from the reference image, specifically depth and normal information. The reference depth $\bm{\hat{I}_{d}}$ and normal $\bm{\hat{I}_{n}}$ are derived using the human foundation model Sapiens [21]. To enforce consistency between the rendered and reference values, we impose depth and normal losses, $\mathcal{L}_{\textnormal{depth}}$ and $\mathcal{L}_{\textnormal{normal}}$. The normal loss is calculated as the mean squared error (MSE) between the reference normal $\bm{\hat{I}_{n}}$ and $\bm{I_{n}}$, i.e., $\mathcal{L}_{\textnormal{normal}}=||\bm{\hat{I}_{n}}-\bm{I_{n}}||_{2}$. For the depth loss, we employ the normalized negative Pearson correlation:
@@
To further refine the geometric details, we adopt DMTet [52], a hybrid SDF-Mesh representation that enables memory-efficient, high-resolution 3D shape reconstruction. The trained NeRF [36] is converted into a mesh, which serves as the geometric initialization for DMTet optimization. During optimization, we compute the reconstruction loss by applying both the MSE loss and the perceptual loss [54] between the rendered normal from the reference view and the reference normal $\bm{\hat{I}_{n}}$. Inspired by HumanNorm [17], we leverage its pretrained normal-adapted and depth-adapted diffusion models for novel view guidance to craft high-quality geometry. Furthermore, the SDF loss $\mathcal{L}_{\textnormal{sdf}}$ is incorporated to regularize the geometry, preventing significant deviation from the initial mesh and avoiding improper shapes. Through geometry initialization and sculpting, we achieve high-fidelity 3D human geometry with plausible poses and intricate geometric details.
Table 1: Quantitative Comparison with State-of-the-art Methods. The evaluation images are selected from in-the-wild images found on the Internet and the CAPE dataset [33]. The best results are highlighted in bold. The second-place results are underlined.
in-the-wild CAPE
[…]
PIFu 26.968 0.035 0.594 26.912 0.028 0.764
GTA 25.060 0.064 0.568 30.376 0.019 0.785
TeCH 25.740 0.053 0.713 27.601 0.025 0.826
SiTH 20.412 0.129 0.608 21.992 0.048 0.815
GeneMAN (Ours) 32.238 0.013 0.730 28.490 0.015 0.838
@@ 4.4 Multi-Space Texture Refinement
where $\lambda^{c}_{ref}$ denotes the weighting parameter of the reference loss, and $\mathcal{L}^{c}_{\textnormal{guid}}$ represents the guidance loss in the coarse texture stage, as defined in Eq. 6.
Multi-View Training. Multi-view training strategies effectively enhance consistency across views [28, 53], but they typically require retraining diffusion models. In contrast, we employ a training-free approach by adding identical Gaussian noise to a batch of images rendered from different views. These images are then concatenated into a single image for inference, ensuring that the resulting multi-view images exhibit consistent textures and colors.
The coarse texture stage prioritizes learning consistent and plausible texture but still lacks realism. To enhance texture realism, we employ pixel-space optimization by minimizing the distance between the rendered multi-view images and their refined counterparts, thereby improving the UV texture map. Akin to DreamGaussian [59], we adopt the image-to-image synthesis framework of SDEdit [35] to generate the refined images. Specifically, we render the coarse image $\bm{I_{\textnormal{corase}}}$ from an arbitrary camera view $c$, perturb it with random Gaussian noise, and apply a multi-step denoising process using the GeneMAN 2D prior based ControlNet [68] $\phi_{2d}(\cdot)$ to obtain the refined image $\bm{I_{\textnormal{fine}}}$:
@@
where $\epsilon(t_{\textnormal{start}})$ represents random noise at timestep $t_{\textnormal{start}}$, and $e$ denotes the conditioned text embeddings. We employ the MSE loss for pixel-wise reconstruction, supplemented by the LPIPS [69] loss to enhance texture details and style similarity. The total loss for pixel space optimization $\mathcal{L}_{\textnormal{fine}}$ is:
@@ 5.1 Implementation Details
Training Details. Our framework is built upon the open-source project ThreeStudio [10]. During the geometry stage, we progressively increase the resolution of NeRF [36] from $256$ to $384$ over $5,000$ steps. We then convert it to an explicit mesh, which serves as the geometry initialization for DMTet [52] at a resolution of $512$. We subsequently optimize DMTet for $3,000$ steps to sculpt fine-grained geometric details. In the texture stage, we perform an initial coarse texture optimization over $10,000$ steps, followed by a refinement of the texture UV map for $1,000$ steps. The full optimization process takes approximately $1.4$ hours. More details are placed in the supplementary.
Testing Details. For both qualitative and quantitative evaluation, we randomly select $50$ samples from the Internet and CAPE [33]. The samples sourced from the Internet include challenging scenarios, such as humans with varying body proportions (head shots, half-body shots, full-body shots), as well as humans in diverse poses, clothing, and with personal belongings. We compare our model with state-of-the-art image-to-3D human reconstruction methods, including PIFu [47], GTA [70], TeCH [18] and SiTH [13]. For each method, we render 120 viewpoints across 360 degrees of the reconstructed results for evaluation. We compute PSNR and LPIPS [69] between the reference view and the rendered front view to evaluate reconstruction quality. Additionally, CLIP similarity [43] between the reference view and $119$ novel views is measured to assess multi-view consistency.
Figure 5: Qualitative GeneMAN Results with Complex Poses.